OpenStudy (anonymous):

At a manufacturing plant, parts are regularly tested. Past records indicate that 2% of all parts tested are found to be defective. In a test batch of 400 parts, how many are expected to be defective?

OpenStudy (anonymous):

2/100 *400=8 parts

OpenStudy (anonymous):

percentage indicates a fraction that is 1/100..... so 2% is 2/100 2% of 400=2/100 *400 =8 so 8 a expected to be defective

OpenStudy (anonymous):

so it's 8